The angle between lines 2x2  -7xy+3y2=0 please say me answer

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
Angle between two lines is given by
 
tan theta = 2 sqrt(h^2 – ab) /(a +b) = 2 sqrt(49/4 – 6) /5 = 2 *5/2 / 5 = 5/5 = 1
 
hence tan theta = 1
theta = 45 degree
 
Hope it helps